# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 2. Jordan type functions for $\mathbb{G}$-modules
Definition 2.15.
We denote by $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y}_{\leq,m})$ the set of continuous maps from $\mathbb{P}V_{r}(\mathbb{G})$ with the Zariski topology to $\mathcal{Y}_{\leq,m}$ with the Alexandrov discrete topology (as in Theorem 1.11). For $F,G\in Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y}_{\leq,m})$, we write $F\leq G$ if and only if $F(\overline{x})\leq G(\overline{x})\in\mathcal{Y}_{\leq,m}$ for all scheme-theoretic points $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$.
The Jordan type function in the following theorem was first formulated in [13, Defn 2.3] and its continuity briefly specified in [13, Prop 2.8.3].
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. If $M$ is a $\mathbb{G}_{(r)}$-module of dimension $m$ and if $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ is a lift of a scheme-theoretic point $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(\overline{x})\to\mathbb{P}V_{r}(\mathbb{G})$, we define
$JT_{\mathbb{G},r,M}(\overline{x})\quad\equiv\quad JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M))$
to be the Jordan type of the $k(x)[u]/u^{p}$-module obtained by restricting $k[V_{r}(\mathbb{G})]\otimes M$ along $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ (equivalent to the Jordan type of $(\mu_{x})_{*}(1\otimes u_{r-1})$ acting on $k(x)\otimes M$).
@@
    The association $\overline{x}\mapsto JT_{\mathbb{G},r,M}(\overline{x})$ determines a continuous map
    $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}_{\leq,m};$
    in other words, $JT_{\mathbb{G}_{(r)},M}(-)\ \in\ Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y}_{\leq,m})$.
@@
    $JT_{\mathbb{G},r,M\oplus N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})+JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{\leq,m+n}.$
@@
    $JT_{\mathbb{G},r,M\otimes N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})\otimes JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{\leq,m\cdot n}.$
@@
    If $f:M\to N$ is a surjective map of $\mathbb{G}_{(r)}$-modules, then $JT_{\mathbb{G},r,N}(\overline{x})\ \leq\ JT_{\mathbb{G},r,M}(\overline{x})$ for any $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$; in other words, $JT_{\mathbb{G},r,N}(-)\leq JT_{\mathbb{G},r,M}(-)$, where the partial ordering of functions is that inherited by the partial ordering of $\mathcal{Y}$.
@@
    Consider a short exact sequence $\mathcal{E}:0\to N\ \to\ M\ \to\ Q\to 0$ of finite dimensional $\mathbb{G}_{(r)}$-modules. Then $\mathcal{E}$ is locally split if and only if $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G},r,N}(-)+JT_{\mathbb{G},r,Q}(-).$
@@
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of finite type over $k$, then
    $JT_{\mathbb{H},r,\phi^{*}(M)}(-):\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathcal{Y}_{\leq,m}$
@@
    $JT_{\mathbb{G},r,M}(-)\circ\phi_{*}:\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathbb{P}V_{r}(\mathbb{G})\ \to\ \mathcal{Y}_{\leq,m}.$
@@
The fact that $JT_{\mathbb{G},r,M}(\overline{x})$ is well defined (namely, independent of the choice of $x\in V_{r}(\mathbb{G})$ projecting to $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$), follows from the equality (2.14.1) of Theorem 2.14.
To prove the asserted continuity of assertion (1) it suffices to prove continuity of $JT_{\mathbb{G},r,M}(-)$ when restricted to each affine open subset of the form $\operatorname{Spec}\nolimits(k[V_{r}(\mathbb{G})][1/h])_{0}\\ \subset\mathbb{P}V_{r}(\mathbb{G})$ where $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$ is a power of some homogeneous generator of $k[V_{r}(\mathbb{G})]$. This follows from the continuity proved in Theorem 1.11 for $A=(k[V_{r}(\mathbb{G})][1/h])_{0}$.
The fact that $JT_{\mathbb{G},r,M\oplus N}(-)=JT_{\mathbb{G},r,M}(-)+JT_{\mathbb{G},r,N}(-)$ is self evident.
Since $(\mu_{x})_{*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ is a map of Hopf algebras (over $k(x)$), the restriction map $\mu_{x}^{*}:mod(k(x)\otimes k\mathbb{G}_{(r)})\to mod(k(x)\otimes k\mathbb{G}_{a(r)})$ commutes with tensor products. Since $\epsilon_{1}$ is the identity, this implies assertion (3) whenever $r=1$. On the other hand, if $x$ is a generic point of $V_{r}(\mathbb{G})$ for any $r>0$, the fact that $M\ \mapsto\ JT_{\mathbb{G},r,M}(x)$ commutes with tensor products is given by [15, Prop4.7].
A surjective map $f:M\to N$ of finite dimensional $\mathbb{G}_{(r)}$-modules induces a surjective map $k(x)\otimes M\ \to\ k(x)\otimes N$ of $k(x)[u]/t^{p}$-modules. Consequently, assertion (4) follows from Corollary 1.4. Assertion (5) follows from the definition of a locally split exact sequence (see [4]) and the fact for any field $K$ and any short exact sequence $0\to N_{K}\to M_{K}\to Q_{K}\to 0$ that the Jordan type of $M_{K}$ equals the sum of the Jordan types of $N_{K}$ and $Q_{K}$ if and only if the short exact sequence splits,
@@
Remark 2.17.
For any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$, $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G},r,M^{*}}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}_{\leq,m}$ where $M^{*}$ is the $k$-linear dual of $M$. Moreover, if $I$ and $J$ restrict to injective $\mathbb{G}_{(r)}$ modules of the same dimension, then $JT_{\mathbb{G},r,I}(-)=JT_{\mathbb{G},r,J}(-)$. Thus, $M\mapsto JT_{\mathbb{G},r,M}(-)$ is far from faithful.
Proposition 2.18.
Consider a linear algebraic group $\mathbb{G}$ and denote by $\mathbb{P}V_{r}(\mathbb{G})(k)/\mathbb{G}(k)$ the quotient of the set of $k$-points of $\mathbb{P}V_{r}(\mathbb{G})$ by the adjoint action of the discrete group $\mathbb{G}(k)$. For any $\mathbb{G}$-module $M$ of dimension $m$, the functor $JT_{\mathbb{G},r,M}(-)$ of Theorem 2.16(1) induces a set map
$JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})(k)/\mathbb{G}(k)\quad\to\quad\mathcal{Y}_{\leq,m}.$
@@
Consider two $k$-points $x,y\in V_{r}(\mathbb{G})$ and $g\in\mathbb{G}(k)$ with $y=g\circ x$ for the adjoint action of $\mathbb{G}(k)$ on $V_{r}(\mathbb{G})$. Let $M^{g}$ denote the $\mathbb{G}$-module with the same underlying $k$-vector space structure as that of $M$ and with $\mathbb{G}$-action as that given by the $\mathbb{G}$-action on $M$ preceded by conjugation by $g$. Observe that action by $g$ determines an isomorphism $g:M\to M^{g}$.
We have a commutative diagram
(2.18.1) ${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$$\scriptstyle{\Theta_{\mathbb{G},r,M^{g}}}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Theta_{\mathbb{G},r,M}}$${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$$\scriptstyle{\chi_{y}^{*}}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi_{x}^{*}}$${k(x)\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${k(y)\otimes M.}$
This implies that the action of $(\mu_{y})_{*}(u_{r-1})=(\chi_{y}^{*}\circ\Theta_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ on $M^{g}$ is isomorphic to the action of $(\mu_{x})_{*}(u_{r-1})=(\chi_{x}^{*}\circ\Theta_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ on $M$. ∎
The following theorem associates to a finite dimensional $\mathbb{G}$-module $M$ the configuration $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\}$ which gives considerably more information about $M$ than does the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}\subset\mathbb{P}V_{r}(\mathbb{G})$ of $M$. The fact that $(V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\subset V_{r}(\mathbb{G})$ is closed is given with a sketch of proof in [13, Prop 2.8.3].
Theorem 2.19.
(see Proposition 2.8.3 of [13]) Let $\mathbb{G}$ be an affine group of finite type over $k$ and let $r$ be a positive integer. Consider a $\mathbb{G}_{(r)}$-module $M$ of dimension $m$. For each $\underline{a}\in\mathcal{Y}_{\leq,m}$, we define the subset
Then $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}};\ \underline{a}\in\mathcal{Y}_{\leq,m}\}$ is a partially ordered configuration of closed subspaces of $\mathbb{P}V_{r}(\mathbb{G})$.
With this notation, the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}$ of [23] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by $\underline{a}\in\mathcal{Y}_{\leq,m}$ which satisfy the condition that $a_{i}\not=0$ for some $i<p$.
The non-maximal support variety $\Gamma(\mathbb{G}_{(r)})_{M}$ of [13] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by the (finitely many) $\underline{a}\in\mathcal{Y}_{\leq,m}$ which satisfy the condition that there exists some $\underline{b}\in\mathcal{Y}_{\leq,m}$ with $\underline{a}\leq\underline{b}$ such that $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is strictly contained in $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$
Proof.
To prove that $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is closed in $\mathbb{P}V_{r}(\mathbb{G})$, it suffices to verify the affine analogue: $V_{r}(\mathbb{G})_{M,\leq\underline{a}}$ is closed in $V_{r}(\mathbb{G})$, where $V_{r}(\mathbb{G})_{M,\leq\underline{a}}\equiv\{x:JT_{\mathbb{G}_{(r)},M}(x)\leq\underline{a}\}$. This follows from Theorem 1.11 with $A=k[V_{r}(\mathbb{G}]$, $M_{A}=A\otimes M$, and $\theta_{A}$ equal to $\Theta_{\mathbb{G},r}$ (of Theorem 2.11).
The identification of the subspaces $\mathbb{P}V_{r}(\mathbb{G})_{M}=\cup_{\underline{a}\not=n[p]}(\mathbb{P}V_{r}(\mathbb{G})_{M,\leq\underline{a}})$ and $\Gamma(\mathbb{G}_{(r)})_{M}$ (as asserted above) inside $\mathbb{P}V_{r}(\mathbb{G})$ follows immediately from the definitions of these subspaces. ∎
We make explicit the following corollary first stated and proved in [13, Cor 2.15]. The configuration $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}};\ \underline{a}\in\mathcal{Y}_{\leq,m}\}$ gives a geometric “picture” of the array of Jordan types of a $\mathbb{G}$-module of dimension $m$. One may view this configuration as giving a geometric interpretation of $JT_{\mathbb{G},r,M}(-)\in Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$.
@@
is locally closed. More precisely, $(\mathbb{P}V_{r}(\mathbb{G}))_{M,=\underline{a}}$ is the complement inside $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ of the union of those $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$ such that $\underline{b}\leq\underline{a}$ and $\underline{b}\not=\underline{a}$,
@@
In [13, Defn 4.1], generalized support varieties $\Gamma^{j}(\mathbb{G})_{M}$ were introduced for finite dimensional $\mathbb{G}$-modules $M$ for an arbitrary group scheme $\mathbb{G}$. For $\mathbb{G}$ infinitesimal, one can recover $\Gamma^{j}(\mathbb{G})_{M}$ as in Theorem 2.19 by replacing $\Theta_{\mathbb{G},r,M}$ in (2.9.1) in Theorem 2.11 by its $j$-th power, then replacing $JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M)$ in Theorem 2.16 by the Jordan type of the $j$-th power of $(\mu_{x})_{*}(u_{r-1})$ acting on $k(x)\otimes M$.

### v2

@@ 2. Jordan type functions for $\mathbb{G}$-modules
Remark 2.19.
For any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$ with $k$-linear dual $M^{*}$,
$JT_{\mathbb{G},r,M}(-)\quad=\quad JT_{\mathbb{G},r,M^{*}}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}_{m}.$
Moreover, if $I$ and $J$ restrict to injective $\mathbb{G}_{(r)}$ modules of the same dimension, then $JT_{\mathbb{G},r,I}(-)=JT_{\mathbb{G},r,J}(-)$. Thus, $M\mapsto JT_{\mathbb{G},r,M}(-)$ is far from faithful.
The next proposition verifies that $JT_{\mathbb{G},r,M}$ is a “function on $\mathbb{G}$-conjugacy classes”. Fix the convention that the action $(-)^{g}:G\to G$ for a group $G$ with $g\in G$ sends $h\in G$ to $g^{-1}\cdot h\cdot g\in G$. This easily leads to the adjoint action of $\mathbb{G}$ on itself (and on its normal subgroups $\mathbb{G}_{(r)}$). Composition of a 1-parameter group $\mathbb{G}_{a(r)}\to\mathbb{G}$ with this adjoint action determines an “adjoint action” of $\mathbb{G}$ on $V_{r}(\mathbb{G})$.
Proposition 2.20.
Consider an affine group scheme $\mathbb{G}$ of finite type over $k$, $r$ a positive integer, $M$ a $\mathbb{G}$-module of dimension $m$, and $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ a scheme-theoretic point of $V_{r}(\mathbb{G})$ corresponding to the 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{k(x)}$. For $g\in\mathbb{G}(k(x))$, denote by $\chi_{x^{g}}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ the the scheme-theoretic point corresponding to the 1-parameter subgroup $(-)^{g}\circ\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{k(x)}$. Then
$JT_{\mathbb{G},r,M}(x^{g})\quad=\quad JT_{\mathbb{G},r,M}(x).$
Proof.
Using base change from $k$ to $k(x)$, we reduce to the notationally simpler case that $k=k(x)$. Let $M^{g}$ denote the $\mathbb{G}$-module with the same underlying vector space structure as that of $M$ and with $h\in\mathbb{G}(k)$ sending $n\in M^{g}$ to $(g^{-1}hg)\cdot n\in M^{g}$ where $\cdot$ refers to the given action of $\mathbb{G}(k)$ on $M$. Observe that the map $g:M^{g}\to M,\ n\mapsto g\cdot n$ is $\mathbb{G}(k)$-equivariant.
This leads to a commutative diagram (viewing group schemes as functors from commutative $k$-algebras to groups)
(2.20.1) ${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes g\bullet}$$\scriptstyle{\Theta_{\mathbb{G},r,M^{g}}}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Theta_{\mathbb{G},r,M}}$${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes g\bullet}$$\scriptstyle{(\mu_{x^{g}})^{*}\otimes 1}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mu_{x})^{*}\otimes 1}$${k[V_{r}(\mathbb{G}_{a})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes g\bullet}$${k[V_{r}(\mathbb{G}_{a})]\otimes M.}$
Consequently, the action of $(\mu_{x^{g}})_{*}(1\otimes u_{r-1})=(\Theta_{\mathbb{G},r,x^{g}})_{*}(1\otimes u_{r-1})$ on $M^{g}$ is isomorphic to the action of $(\mu_{x})_{*}(1\otimes u_{r-1})=(\Theta_{\mathbb{G},r,x})_{*}(1\otimes u_{r-1})$ on $M$, where $\Theta_{\mathbb{G},r,x^{g}}$ and $\Theta_{\mathbb{G},r,x}$ are the specializations of $\Theta_{\mathbb{G},r}$ at $x^{g}$ and $x$. ∎
In the theorem below, we equip (the scheme-theoretic points of) $\mathbb{P}V_{r}(\mathbb{G})$ with the Zariski topology and equip $\mathcal{Y}_{m}$ with the Alexandrov discrete topology (see Theorem 1.12). We refer to [15, Prop 2.8.3] for an earlier discussion of properties of the Jordan type function.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer.
@@
    For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, the function $JT_{\mathbb{G},r,M}(-)$ of (2.16.1) induces a continuous map
    $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}$
    sending a scheme-theoretic point $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $JT_{\mathbb{G},r,M}(x)$ for any choice of scheme-theoretic point $x\in V_{r}(\mathbb{G})$ projecting to $\overline{x}$.
  2. (2)
    The finite dimensional $\mathbb{G}_{(r)}$-module $M$ is injective if and only if the dimension $m$ of $M$ is divisible by $p$ and $JT_{\mathbb{G},r,M}(x)\ =\ \frac{m}{p}\cdot[p]$ for all $x\in\mathbb{P}V_{r}(\mathbb{G})$.
@@
    $JT_{\mathbb{G},r,M\oplus N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})+JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{m+n}.$
@@
    $JT_{\mathbb{G},r,M\otimes N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})\otimes JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{m\cdot n}.$
@@
    If $f:M\to N$ is a surjective map of finite dimensional $\mathbb{G}_{(r)}$-modules, then $JT_{\mathbb{G},r,N}(\overline{x})\ \leq\ JT_{\mathbb{G},r,M}(\overline{x})$ for any $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$; in other words, $JT_{\mathbb{G},r,N}(-)\leq JT_{\mathbb{G},r,M}(-)$, where the partial ordering of functions is that inherited by the partial ordering of $\mathcal{Y}$.
@@
    Consider a short exact sequence $\mathcal{E}:0\to N\ \to\ M\ \to\ Q\to 0$ of finite dimensional $\mathbb{G}_{(r)}$-modules. Then $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G},r,N}(-)+JT_{\mathbb{G},r,Q}(-)$ if and only if the short exact sequence $\mu_{x}^{*}(\mathcal{E})$ of $\mathbb{G}_{k(x)}$-modules splits for all $x\in V_{r}(\mathbb{G})$.
@@
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of finite type over $k$, then for any finite dimensional $\mathbb{G}$-module $M$
    $JT_{\mathbb{H},r,\phi^{*}(M)}(-):\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathcal{Y}$
@@
    $JT_{\mathbb{G},r,M}(-)\circ\phi_{*}:\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathbb{P}V_{r}(\mathbb{G})\ \to\ \mathcal{Y}.$
@@
The fact that $JT_{\mathbb{G},r,M}(\overline{x})$ is well defined (namely, independent of the choice of $x\in V_{r}(\mathbb{G})$ projecting to $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$), follows from the equality (2.18.1) of Theorem 2.18. To prove the asserted continuity of assertion (1) it suffices to prove continuity of $JT_{\mathbb{G},r,M}(-)$ when restricted to each affine open subset of the form $\operatorname{Spec}\nolimits(k[V_{r}(\mathbb{G})][1/h])_{0}\subset\mathbb{P}V_{r}(\mathbb{G})$, where $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$ is an irreducible homogeneous generator of $k[V_{r}(\mathbb{G})]$ raised to the power $p^{r-1}/deg(h)$. This follows from the continuity proved in Theorem 1.12 for $A=(k[V_{r}(\mathbb{G})][1/h])_{0}$.
Assertion (2) follows from the detection theorem for nilpotence for $\mathbb{G}_{(r)}$-modules given in [27, Thm 4.3, Cor 4.4].
The fact that $JT_{\mathbb{G},r,M\oplus N}(-)=JT_{\mathbb{G},r,M}(-)+JT_{\mathbb{G},r,N}(-)$ is self evident. Since $(\mu_{x})_{*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ is a map of Hopf algebras (over $k(x)$), the restriction map $\mu_{x}^{*}:mod(k(x)\otimes k\mathbb{G}_{(r)})\to mod(k(x)\otimes k\mathbb{G}_{a(r)})$ commutes with tensor products. Since $\epsilon_{1}$ is the identity, this implies assertion (4) whenever $r=1$. On the other hand, if $x$ is a generic point of $V_{r}(\mathbb{G})$ for any $r>0$, the fact that $M\ \mapsto\ JT_{\mathbb{G},r,M}(x)$ commutes with tensor products is given by [17, Prop 4.7].
A surjective map $f:M\to N$ of finite dimensional $\mathbb{G}_{(r)}$-modules induces a surjective map $k(x)\otimes M\ \to\ k(x)\otimes N$ of $k(x)[u]/t^{p}$-modules. Consequently, assertion (5) follows from Corollary 1.4. Assertion (6) follows from the definition of a locally split exact sequence (see [6]) and the fact for any field $K$ and any short exact sequence $0\to N_{K}\to M_{K}\to Q_{K}\to 0$ that the Jordan type of $M_{K}$ equals the sum of the Jordan types of $N_{K}$ and $Q_{K}$ if and only if the short exact sequence splits.
@@
The following corollary is rephrasing of some of the properties of Theorem 2.21.
Corollary 2.22.
We denote by $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ the set of continuous maps from $\mathbb{P}V_{r}(\mathbb{G})$ with the Zariski topology to $\mathcal{Y}$ with the Alexandrov discrete topology. Give (the scheme-theoretic points of ) $\mathbb{P}V_{r}(\mathbb{G})$ the partial ordering of specialization and give $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ the partial ordering inherited from that of $\mathcal{Y}$.
Then sending a finite dimension $\mathbb{G}_{(r)}$-module $M$ to $JT_{\mathbb{G},r,M}(-)$ determines
$JT_{\mathbb{G},r}:mod(\mathbb{G}_{(r)})\quad\to\quad Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$
commuting with sums and the partial orderings.
The following theorem associates to a finite dimensional $\mathbb{G}$-module $M$ the configuration $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\}$ which gives considerably more information about $M$ than does the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}\subset\mathbb{P}V_{r}(\mathbb{G})$ of $M$. The fact that $(V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\subset V_{r}(\mathbb{G})$ is closed is given with a sketch of proof in [15, Prop 2.8.3].
Theorem 2.23.
(see Proposition 2.8.3 of [15]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. Consider a $\mathbb{G}_{(r)}$-module $M$ of dimension $m$. For each $\underline{a}\in\mathcal{Y}_{m}$, we define the subset
Then $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}};\ \underline{a}\in\mathcal{Y}_{m}\}$ is a partially ordered configuration of closed subspaces of $\mathbb{P}V_{r}(\mathbb{G})$.
With this notation, the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}$ of [27] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by $\underline{a}\in\mathcal{Y}_{m}$ which satisfy the condition that $a_{i}\not=0$ for some $i<p$. (See Theorem 2.6.)
The non-maximal support variety $\Gamma(\mathbb{G}_{(r)})_{M}$ of [15] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by those $\underline{a}\in\mathcal{Y}_{m}$ which satisfy the condition that there exists some $\underline{b}\in\mathcal{Y}_{m}$ with $\underline{a}\leq\underline{b}$ such that $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is strictly contained in $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$
@@
The continuity of $JT_{\mathbb{G},r,M}(-)$ together with the evident fact for any $\underline{a}\in\mathcal{Y}$ that $\{\underline{c}\in\mathcal{Y}:\ \underline{c}\leq a\}$ is closed in $\mathcal{Y}$ implies $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is closed in $\mathbb{P}V_{r}(\mathbb{G})$. The identification of the subspaces $\mathbb{P}V_{r}(\mathbb{G})_{M}=\cup_{\underline{a}\not=n[p]}(\mathbb{P}V_{r}(\mathbb{G})_{M,\leq\underline{a}})$ and $\Gamma(\mathbb{G}_{(r)})_{M}$ (as asserted above) inside $\mathbb{P}V_{r}(\mathbb{G})$ follows immediately from the definitions of these subspaces. ∎
We make explicit the following corollary first stated and proved in [15, Cor 2.15].
@@
is locally closed, the union of those $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$ such that $\underline{b}\leq\underline{a}$ and $\underline{b}\not=\underline{a}$,
@@
We can interpret Corollary 2.24 as asserted that the Jordan type function $JT_{\mathbb{G},r,M}(-)$ for a finite dimensional $\mathbb{G}_{(r)}$-module $M$ determines a stratification of $\mathbb{P}V_{r}(\mathbb{G})$ with locally closed strata, each colored by a Young diagram such that if one stratum is strictly contained in the closure of another stratum then the Young diagram of the smaller stratum is less than the Young diagram of the bigger stratum.
Remark 2.26.
In [15, Defn 4.1], generalized support varieties $\Gamma^{j}(\mathbb{G})_{M}$ were introduced for finite dimensional $\mathbb{G}$-modules $M$ for an arbitrary finite group scheme $\mathbb{G}$ over $k$. Namely, $\Gamma^{j}(\mathbb{G})_{M}$ is the subspace of those $\pi$-points $\alpha_{K}]\in\Pi(\mathbb{G})$ with the property that the rank of $\alpha_{K}(t^{j})$ on $M_{K}$ is not maximal. For $\mathbb{G}$ infinitesimal, one can recover $\Gamma^{j}(\mathbb{G})_{M}$ as in Theorem 2.23 by replacing $\Theta_{\mathbb{G},r,M}$ in Theorem 2.15 by its $j$-th power, then replacing $JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M)$ in Theorem 2.21 by the Jordan type of the $j$-th power of $(\mu_{x})_{*}(u_{r-1})$ acting on $k(x)\otimes M$.